Life Cycle of High-Density Polyethene

Andrew Lorimer

May 2018

Abstract

High-density polyethene (HDPE) is a practical and economical plastic used for pipes, bottles, containers, construction materials, and several other uses. It is inherently UV stable and chemically quite resistant - it is almost impossible to damage with acids or bases. Further, HDPE is easily recycled, and has a resin identification code of 2. In 2015, the HDPE market was worth about $80 billion (AUD).

1 Invention

Figure 1: amorphous vs crystalline polymers High-density polyethene (polyethylene) is a multiple of the basic ethene molecule (H_{2}CCH_{2}). Whilst ethene occurs naturally in plants, polyethene must be synthesised. In 1898, a German chemist Hans von Pechmann created a compound called polymethylene, which was the basis for the invention of polyethylene. In 1933, Fawcett and Gibson synthesised polyethylene using high pressure and a small proportion of oxygen at the British Imperial Chemical Industries. This is illustrated below:

(H_{2}CCH_{2})n ⟶_{< 10ppm O2}^{≈0.2 GPa} (CH_{2}CH_{2})n

However, this high-pressure method can only be used to synthesise low-density polyethene because the high pressure causes a lot of random chain branching (i.e. the molecules are non-linear). Linear molecules used in HDPE are stronger due to their larger intermolecular forces resulting from their surface area, giving them a more amorphous (shapeless) structure – see fig. 1.

During the 1940s-50s, two chemists independently discovered methods of synthesising HDPE using catalysts (Karl Zielger used alkylaluminium and titanium, and Phillips Petroleum used chromium trioxide).

( H

2

CCH

2

) n

Al-based catalyst

⟶

≈0.2 GPa

( CH

2

CH

2

)

n

Later in the 1950s, HDPE was produced industrially using both methods, primarily by DuPont, and was used for internal shielding in cables, infrared lenses, and generally replaced PVC in many applications.

In the 1970s, HDPE synthesis was advanced by using magnesium chloride as a catalyst (

References